OpenStudy (anonymous):

Mario opens a store credit card when he purchases new carpeting. The interest rate is 20.45%. He charges $3,240 to the card, and can pay $1,150 per month. What will the total cost of his purchase be?

OpenStudy (tkhunny):

Month 0: 3240 Month 1: 3240(1 + 0.2045/12) - 1150 = ? Month 2: ?(1 + 0.2045/12) - 1150 = ?? Month 3: ??(1 + 0.2045/12) - 1150 = ??? Did we pay it off, yet?